OpenStudy (flowey15):

An office building casts a 90 foot shadow. At the same time of day, a 5 foot woman standing near the building casts an 8 foot shadow. What is the approximate height of the building to the nearest foot? A) 40 feet B) 49 feet C) 56 feet D) 63 feet

OpenStudy (anonymous):

If a five foot woman makes an 8 foot shadow. That means every 8 feet of shadow is 5 feet of actual height. So we do 90/8 * 5 = ?

OpenStudy (flowey15):

which is 56.25?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Solve the following for x: \[\frac{5}{8}=\frac{x}{90} \]